ristar Production Company began operations on September 1, 2013. Listed below are a number of transactions that occurred during its first four months of operations. (FV of $1, PV of $1, FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided.) 1. On September 1, the company acquired five acres of land with a building that will be used as a warehouse. Tristar paid $200,000 in cash for the property. According to appraisals, the land had a fair value of $134,200 and the building had a fair value of $85,800. 2. On September 1, Tristar signed a $50,000 noninterest-bearing note to purchase equipment. The $50,000 payment is due on September 1, 2014. Assume that 8% is a reasonable interest rate. 3. On September 15, a truck was donated to the corporation. Similar trucks were selling for $3,500. 4. On September 18, the company paid its lawyer $8,000 for organizing the corporation. 5. On October 10, Tristar purchased machinery for cash. The purchase price was $25,000 and $1,000 in freight charges also were paid. 6. On December 2, Tristar acquired various items of office equipment. The company was short of cash and could not pay the $6,500 normal cash price. The supplier agreed to accept 200 shares of the company's nopar common stock in exchange for the equipment. The fair value of the stock is not readily determinable. 7. On December 10, the company acquired a tract of land at a cost of $30,000. It paid $2,500 down and signed a 10% note with both principal and interest due in one year. Ten percent is an appropriate rate of interest for this note. Required: Prepare journal entries to record each of the above transactions.
Answer #1

Solution

Journal entries

Serial

No.

Date

account

Titles

Debit credit
1. Sep 1 land* 122000
Building** 78000
Cash 200000
2. Sep 1 equipment*** 46296
Discount on notes payable 3704
Notes payable 50000
3. Sep15 Truck 3500
Revenue:donation of asset 3500
4. Sep18 organization cost expenses 8000
Cash 8000
5. Oct10 Machinery 26000
Cash 26000
6. Dec2 office equipment 6500
Common stock 6500
7. Dec 10 land 30000

Cash

2500
Notes payable 27500

*land=134200/220000×2000000=122000

**Building=85800/220000×200000=78000

***PV=50000(1.08)^-1=46296